OpenStudy (anonymous):

Find the absolute minimum and absolute maximum values of f on the given interval. f(x) = x − ln 8x, 1 2 , 2

OpenStudy (freckles):

First step: Find derivative Second step: Find critical numbers that occur between the endpoints. Third Step: Plug in critical numbers that are between the endpoints and also plug in endpoints into f.

OpenStudy (freckles):

Can find the derivative?

OpenStudy (freckles):

the endpoints are given

OpenStudy (freckles):

I will walk you through it. But I would like you to first find f'(x)

OpenStudy (freckles):

(x)'=? (ln(8x))'=?

OpenStudy (freckles):

And I think you mean the interval is [1/2,2]

OpenStudy (freckles):

1/2 is the left endpoint 2 is the right endpoint

OpenStudy (freckles):

Do you know (x)'=?

OpenStudy (freckles):

What is the slope of the line y=x?

OpenStudy (freckles):

\[\frac{d}{dx}(\ln(8x))=\frac{d}{dx}\ln(8)+\frac{d}{dx}\ln(x)=?\]

OpenStudy (freckles):

well it looks like you know the derivative of x but not the derivative of constant or ln(x)

OpenStudy (freckles):

\[\frac{d}{dx}c=0 \text{ where c is a constant } \\ \frac{d}{dx}\ln(x)=\frac{1}{x}\]

OpenStudy (freckles):

\[y'=1-\frac{1}{x}\] Find the critical numbers now.

OpenStudy (freckles):

You need to find when y'=0 and when y' dne (we only care about the values that make y' dne that actually exist in the domain of the original function)

OpenStudy (freckles):

So I find it is always easiest to combine fractions y'=(x-1)/x

OpenStudy (freckles):

no you find when y'=0 and when y' dne

OpenStudy (freckles):

We have a fraction for y' All you have to do to find when y'=0 is to find when the top is equal to 0 To find when y' dne you find when the bottom is 0 (again here we only care about numbers that actually happen in our domain)

OpenStudy (freckles):

no

OpenStudy (freckles):

We are simply finding the critical numbers right now

OpenStudy (freckles):

critical numbers could be min or max or maybe not even either

OpenStudy (freckles):

Right. We also need to consider the critical numbers that we have yet to find.

OpenStudy (freckles):

Let me know when you have the equations and I will check to see if you have the right critical numbers.

OpenStudy (freckles):

y'=0 when top=0 y' dne when bottom=0 and you have y'=(x-1)/x

OpenStudy (freckles):

Just summarizing what we have already said

OpenStudy (freckles):

@garyforest ... have you set the top equal to 0 yet?

OpenStudy (freckles):

I don't mean to be impatient but you should have already been able to solve x-1=0

OpenStudy (freckles):

Right.

OpenStudy (freckles):

so x=1 is a critical number and it is a critical number that occurs in between the endpoints y' dne at x=0 but we don't care about x=0 anyways because ln(0) doesn't even exist so x=0 doesn't even occur in our domain also it doesn't even occur in the given interval

OpenStudy (freckles):

so check x=1/2 , check x=1 , check x=2

OpenStudy (freckles):

\[f(x)=x-\ln(8x)\]

OpenStudy (freckles):

plug each number into your initial

OpenStudy (freckles):

and the smallest number will be your absolute min the largest number will be your absolute max

OpenStudy (freckles):

I will leave this part to you because I must go now